# Soft Masked Transformer for Point Cloud Processing with Skip Attention-Based Upsampling

### v1

@@ III-B2 Enhanced Position Encoding
Figure 1: Comparison of the attention, position encoding in Transformers. (a) The vector attention with position encoding bias in Point Transformer, see Eq.(1). (b) The vector attention with position encoding multiplier in Point Transformer V2, see Eq.(2). (c) The vector attention with soft mask and enhanced position encoding bias in our proposed SMTransformer, see Eq.(3).
@@
Residual connections are instrumental in training deep neural networks, facilitating gradient flow during backpropagation. Therefore, we combine the Soft Masked Transformer with residual connections to construct a transformer block. As illustrated in Fig. 4.(b), the Soft Masked Transformer Block (SMTB) can be expressed as,
$f_{i}$ $={\rm Linear}(f_{in1}),$ (8)
$\mathcal{G}_{i}$ $={\rm SMTransformer}(f_{i}),$
$f_{out1}$ $={\rm Linear}(\mathcal{G}_{i}\oplus f_{i}),$
@@ III-C Skip Attention-based Up-sampling Block (SAUB)
To facilitate deep communication between features over various resolution points, we introduce a skip attention-based up-sampling block that combines conventional unpooling with a learnable unit to learn and refine contextual information between features from the encoding and decoding layers across different resolutions.
As illustrated in Fig. 4.(c), given the skip feature I $f_{h}\in\mathbb{R}^{{M\times C_{h}}}$ and skip feature II $f_{l}\in\mathbb{R}^{{m\times C_{l}}}$ from the two adjacent encoding layers over the low and high-resolution points, respectively. To build the communication between the different resolution point features, we first balance their feature dimension and point resolution,
$f_{mid}={\rm Gridup}\big{(}{\rm Linear}(f_{in2}\oslash f_{l})\big{)},$ (9)
@@
$\mathcal{G}_{i}^{sa}=\sum_{j=1}^{K}\mathcal{A}\big{(}(K_{ij}\ominus Q_{i})\oplus{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\delta(\cdot)}\big{)}\odot\big{(}V_{ij}\oplus{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\delta(\cdot)}\big{)},$ (11)
@@
To compare the differences between SAUB and classical upsampling block ( i.e. Transition up[22]), we illustrate their diagrams in Fig. 2. The key differences lie in two aspects:
i) The classical upsampling block only learns the connection between features from the encoding and decoding layers over the same resolution points. In contrast, our proposed SAUB can learn the connection between features from the encoding and decoding layers of different resolution points.
ii) The classical upsampling block uses a simple skip connection to learn the contextual information between the features from the encoding and decoding layers. However, we propose skip attention to refine the contextual information between the features.
Figure 2: Comparison of (a) Transition up and (b) Skip attention-based up-sampling. ‘interpo.’ stands for the interpolation operation. ‘grid up’ stands for the grid-based unpooling. ‘SA’ donates the skip attention.
III-D Shared Position Encoding
In conventional point transformer networks, several transformer blocks typically utilize different position encoding information within the same encoding or decoding layer across resolution points. We refer to this practice as unshared position encoding. Intuitively, points within the same position in the point cloud should have the same position information. Building on this intuition, we propose the shared position encoding strategy. Under this strategy, transformer blocks within the same encoding or decoding layer share identical position encoding information across resolution points. This approach enhances network robustness and efficiency, particularly in large-scale scene processing. Their comparison is illustrated in Fig. 3. The robustness and efficiency experiments are provided in Section IV-D.
Figure 3: (a) Unshared position encoding: Various transformer blocks (coloured in grey) within the same encoding or decoding layer (coloured in yellow), operating over the same resolution point cloud, require different position information (coloured in blue). (b) Shared position encoding: Various transformer blocks within the encoding and decoding layers share position information over the same resolution point cloud.
@@ III-E Network Architecture
We employ a U-Net-like architecture comprising five encoding and decoding layers with skip connections for the semantic segmentation task. The first encoding and decoding layers consist of one MLP and SMTransformer block. The subsequent encoding layers incorporate one Grid Pooling layer [22] followed by several SMTransformer blocks. The number of SMTransformer blocks in the five encoding layers is [1, 2, 2, 6, 2]. The subsequent decoding layers consist of one skip attention-based up-sampling block and one SMTransformer block. We set the feature dimensions $C$ as [32, 64, 128, 256, 512] for the 5 encoding and decoding layers. At the network’s end, we append an MLP to predict the final point-wise labels. The network architecture for semantic segmentation is illustrated in Fig. 4.(a).
For the classification task, we utilize the basic PointNext [47] as the backbone and replace one SMTransformer block with an MLPs block to form the new network architecture for classification. Further details regarding the configurations of the segmentation and classification networks are in Section IV.
@@ IV Experimental Results
We evaluate our network on three tasks namely, semantic segmentation of indoor scenes, semantic segmentation of outdoor scenes, and shape classification. We also perform detailed ablation studies to demonstrate the effectiveness and robustness of the proposed Soft Mask Transformer Block, Skip Attention-based up-sampling Block and the Shared Position Encoding strategy.
TABLE I: Semantic segmentation results on the S3DIS dataset Area-5. We report the mean class-wise Intersection over Union (mIoU), mean class-wise accuracy (mAcc), and overall accuracy (OA). The best result is highlighted in bold, and the second best is underlined.
Year Methods mIoU mAcc OA ceil. floor wall beam column window door chair table bookcase sofa board clut.
2017 CVPR PointNet[24] 41.09 48.98 88.80 97.33 69.80 0.05 3.92 46.26 10.76 52.61 58.93 40.28 5.85 26.38 33.22
2018 NIPS PointCNN[36] 57.26 63.86 85.9 92.3 98.2 79.4 0.0 17.6 22.8 62.1 74.4 80.6 31.7 66.7 62.1 56.7
2019 ICCV KPConv[10] 67.1 72.8 92.8 97.3 82.4 0.0 23.9 58.0 69.0 91.0 81.5 75.3 75.4 66.7 58.9
2020 PAMI SPH3D-GCN[48] 59.5 65.9 93.3 97.1 81.1 0.0 33.2 45.8 43.8 79.7 86.9 33.2 71.5 54.1 53.7
2020 CVPR PointANSL[21] 62.6 68.5 87.7 94.3 98.4 79.1 0.0 26.7 55.2 66.2 86.8 83.3 68.3 47.6 56.4 52.1
2020 CVPR SegGCN[15] 63.6 70.4 93.7 98.6 80.6 0.0 28.5 42.6 74.5 80.9 88.7 69.0 71.3 44.4 54.3
2021 CVPR PAConv[49] 66.6 73.0 94.5 98.6 82.4 0.0 26.4 58.0 60.0 89.7 80.4 74.3 69.8 73.5 57.7
2021 CVPR BAAF-Net[50] 65.4 73.1 88.9 92.9 97.9 82.3 0.0 23.1 65.5 64.9 87.5 78.5 70.7 61.4 68.7 57.2
2021 ICCV Point Transformer[23] 70.4 76.5 90.8 94.0 98.5 86.3 0.0 38.0 63.4 74.3 82.4 89.1 80.2 74.3 76.0 59.3
2022 CVPR CBL[51] 69.4 75.2 90.6 93.9 98.4 84.2 0.0 37.0 57.7 71.9 81.8 91.7 75.6 77.8 69.1 62.9
2022 CVPR RepSurf-U[27] 68.9 76.0 90.2
2022 CVPR Stratified Transformer[41] 72.0 78.1 91.5
2022 ECCV PointMixer[52] 71.4 77.4 94.2 98.2 86.0 0.0 43.8 62.1 78.5 82.2 90.8 79.8 73.9 78.5 59.4
2022 NIPS PointNeXt[47] 71.1 77.2 91.0 94.2 98.5 84.4 0.0 37.7 59.3 74.0 91.6 83.1 77.2 77.4 78.8 60.6
2022 NIPS PointTransformerV2[22] 71.6 77.9 91.1
2023 TCSVT LCPFormer[53] 70.2 76.8 90.8
2023 TCSVT SAKS[54] 68.8 74.0 90.8 95.2 98.6 84.1 0.0 27.5 58.5 75.1 80.4 90.8 69.0 77.0 73.5 62.1
2023 TNNLS PicassoNet++[55] 71.0 77.2 91.3 94.4 98.4 87.5 0.0 46.9 63.7 75.5 81.4 90.3 71.3 76.2 76.7 61.1
2023 CVPR Point Vector[29] 72.3 78.1 91.0 95.1 98.6 85.1 0.0 41.4 60.8 76.7 92.1 84.4 77.2 82.0 85.1 61.4
SMTransformer(ours) 73.4 78.9 91.8 95.2 98.7 87.7 0.0 45.8 64.8 75.2 85.2 92.7 86.7 76.8 83.5 62.4
TABLE II: Semantic segmentation results on S3DIS with 6-fold cross validation.
Year Methods mIoU mAcc OA ceil. floor wall beam column window door chair table bookcase sofa board clut.
2017 CVPR PointNet[24] 47.6 66.2 78.6 88.0 88.7 69.3 42.4 23.1 47.5 51.6 42.0 54.1 38.2 9.6 29.4 35.2
2018 NIPS PointCNN[36] 65.4 75.6 88.1 94.8 97.3 75.8 63.3 51.7 58.4 57.2 69.1 71.6 61.2 39.1 52.2 58.6
2019 CVPR PointWeb[30] 66.7 76.2 87.3 93.5 94.2 80.8 52.4 41.3 64.9 68.1 67.1 71.4 62.7 50.3 62.2 58.5
2019 ICCV KPConv[10] 70.6 79.1 93.6 92.4 83.1 63.9 54.3 66.1 76.6 64.0 57.8 74.9 69.3 61.3 60.3
2020 PAMI SPH3D-GCN[48] 68.9 77.9 88.6 93.3 96.2 81.9 58.6 55.9 55.9 71.7 82.4 72.1 64.5 48.5 54.8 60.4
2020 CVPR PointANSL[21] 68.7 79.0 88.8 95.3 97.9 81.9 47.0 48.0 67.3 70.5 77.8 71.3 60.4 50.7 63.0 62.8
2020 CVPR RandLA-Net[42] 70.0 82.0 88.0 93.1 96.1 80.6 62.4 48.0 64.4 69.4 76.4 69.4 64.2 60.0 65.9 60.1
2021 CVPR PAConv[49] 69.3 78.7 94.3 93.5 82.8 56.9 45.7 65.2 74.9 59.7 74.6 67.4 61.8 65.8 58.4
2021 CVPR SCF-Net[56] 71.6 82.7 88.4 93.3 96.4 80.9 64.9 47.4 64.5 70.1 81.6 71.4 64.4 67.2 67.5 60.9
2021 CVPR BAAF-Net[50] 72.2 83.1 88.9 93.3 96.8 81.6 61.9 49.5 65.4 73.3 83.7 72.0 64.3 67.5 67.0 62.4
2021 ICCV Point Transformer [23] 73.5 81.9 90.2 94.3 97.5 84.7 55.6 58.1 66.1 78.2 74.1 77.6 71.2 67.3 65.7 64.8
2022 NIPS PointNeXt[47] 74.9 83.0 90.3
2022 CVPR RepSurf-U[27] 74.3 82.6 90.8
2022 CVPR CBL[51] 73.1 79.4 89.6 94.1 94.2 85.5 50.4 58.8 70.3 78.3 75.0 75.7 74.0 71.8 60.0 62.4
2023 CVPR Point Vector[29] 78.4 86.1 91.9
SMTransformer(ours) 79.0 86.9 91.9 97.4 98.3 89.4 68.0 66.1 70.4 78.4 82.6 84.0 78.5 72.2 73.2 68.5
@@
TABLE III: Semantic segmentation results (mIoU) on ScanNetV2 validation and test set.
Year Methods Input Val(%) Test(%)
2018 NIPS PointNet++[25] point 55.7 53.5
2018 CVPR SparseConvNet[57] voxel 72.5 69.3
2019 CVPR PointConv[13] point 66.6 61.0
2020 CVPR PointANSL[21] point 63.5 66.6
2019 ICCV MVPNet[58] point 66.4
2019 ICCV KPConv[10] point 69.2 68.6
2019 3DV JointPointBased[59] point 69.2 63.4
2019 CVPR MinkowskiNet[60] voxel 72.2 73.6
2022 CVPR RepSurf-U[27] point 70.0
2022 CVPR Stratified Transformer[41] point 74.3 73.7
2021 CVPR PointTransformer[23] point 70.6
2022 CVPR FastPointTransformer[61] voxel 72.0
2022 NIPS PointTransformerV2[22] point 75.4 75.2
2023 TNNLS PicassoNet++[55] mesh 69.2
SMTransformer (ours) point 75.9 75.7
@@ IV-A Indoor Semantic Segmentation
Datasets: We evaluate our network on two large-scale indoor scene datasets, namely S3DIS [62] and ScanNetV2 [63]. The S3DIS dataset consists of RGB-D point clouds annotated point-wise with 13 classes. It encompasses 271 rooms from 6 large-scale indoor scenes, totalling 6020 square meters. We utilize 6-dimensional point features, including 3-dimensional normalized colour and 3-dimensional normalized location. For evaluation, we conduct a 6-fold cross-validation on S3DIS and focus more extensively on comparisons using Area 5 as the test set, which is distinct from the other areas and not within the same building.
@@
On ScanNet, we set the voxel size as 2cm and the maximum number of voxels to 100,000. We use the Adam optimizer, where the weight decay is set as 0.02. The base learning rate is set as 0.02 and the learning rate is scheduled by the MultiStepLR every 40 epochs. We train and test the model with batch size 24 on 4 GPUs. We adopt random rotation, random scaling, random flip, elastic distortion, chromatic contrast, chromatic translation, chromatic jitter and hue saturation translation to augment training data. The grid size is set as [0.04, 0.12, 0.36, 1.08]cm, and the number of neighbour points is set as 16. During the test, the network uses the test time augmentation, following the Point Transformer V2[22] and Stratified Transformer[41].
Results: We compare our method with the recent state-of-the-art on S3DIS dataset, using three metrics i.e. mean class-wise intersection over union (mIoU), mean overall accuracy (mAcc) and overall accuracy (OA). Results are reported in Table I. Our network demonstrates superior performance on all three metrics i.e. 73.4% mIoU, 78.9% mAcc and 91.8% OA. It achieves the top 2 results on 9 out of 13 classes including ceiling, floor, wall, column, window, table, bookcase, board, clutter. Notably, the segmentation performance of bookcase class exceeds the second-best method by 6.5% mIoU. Compared to the previous state-of-the-art point transformers (e.g. Point Transformer V2 and Stratified Transformer), our network outperforms them by 1.8% and 1.4% in terms of mIoU, respectively. Compared to the MLP-based method (e.g. Point Vector), the performance of our method exceeds it by 1.1% mIoU. Compared to the LCPFormer and SAKS, our method outperforms them by a large margin on all metrics. Fig. 5. shows visualizations of our results on S3DIS area 5 in comparison to Point Transformer V2. We can see that our method is more robust to the object boundaries.
Table II shows results with the 6-fold validation setting on the S3DIS dataset. Our method again achieves state-of-the-art results of 79.0% mIoU, 86.9% mAcc and 91.9% OA. It achieves the best results on 12 out of 13 classes including ceiling, floor, wall, beam, column, window, door, table, bookcase, sofa, board and clutter.
The ScanNetV2 validation and test set results are illustrated in Table III. Compared to the Stratified Transformer, our method exhibits a substantial improvement of +1.6% mIoU and +2.2% mIoU on validation and test sets, respectively. Against the Point Transformer V2, our method delivers enhanced performance with an improvement of (+0.5%, +0.5% in terms of mIoU) on the validation and test sets, respectively.
TABLE IV: Semantic segmentation results on the SemanticKITTI test set. ‘*’ means the network is pre-trained on other datasets.
@@
motorcycle
truck
other-vehicle
person
bicyclist
motorcyclist
road
parking
sidewalk
other-ground
building
fence
vegetation
trunk
terrain
pole
traffic-sign
2017 NIPS Pointnet++[25] 20.1 53.7 1.9 0.2 0.9 0.2 0.9 1.0 0.0 72.0 18.7 41.8 5.6 62.3 16.9 46.5 13.8 30.0 6.0 8.9
2018 ICRA SqueezeSeg[64] 30.8 68.3 18.1 5.1 4.1 4.8 16.5 17.3 1.2 84.9 28.4 54.7 4.6 61.5 29.2 59.6 25.5 54.7 11.2 36.3
2019 ICRA SqueezeSegV2[65] 39.6 82.7 21.0 22.6 14.5 15.9 20.2 24.3 2.9 88.5 42.4 65.5 18.7 73.8 41.0 68.5 36.9 58.9 12.9 41.0
2019 IROS RangNet++[66] 52.2 91.4 25.7 34.4 25.7 23.0 38.3 38.8 4.8 91.8 65.0 75.2 27.8 87.4 58.6 80.5 55.1 64.6 47.9 55.9
2020 CVPR PolarNet[67] 54.3 93.8 40.3 30.1 22.9 28.5 43.2 40.2 5.6 90.8 61.7 74.4 21.7 90.0 61.3 84.0 65.5 67.8 51.8 57.5
2021 CVPR Cylinder3D[68] 68.9 97.1 67.6 63.8 50.8 58.5 73.7 69.2 48.0 92.2 65.0 77.0 32.3 90.7 66.5 85.6 72.5 69.8 62.4 66.2
2021 CVPR (AF)${}^{2}$-S3Net[69] 69.7 94.5 65.4 86.8 39.2 41.1 80.7 80.4 74.3 91.3 68.8 72.5 53.5 87.9 63.2 70.2 68.5 53.7 61.5 71.0
2023 ICCV RangFormer*[73] 73.3 96.7 69.4 73.7 59.9 66.2 78.1 75.9 58.1 92.4 73.0 78.8 42.4 92.3 70.1 86.6 73.3 72.8 66.4 66.6
IV-B Outdoor Semantic Segmentation
Datasets: We conduct experiments on two popular datasets: SemanticKITTI[74] and SWAN [72] dataset. The SemanticKITTI provides 22 sequence point clouds consisting of 43,552 frames. Adhering to standard practice, we employ sequences 0 to 10 (excluding 8) for training, use sequence 8 for validation and sequences 11 to 21 for testing. The labels for the test set are exclusively available to the online server, necessitating result submissions for remote evaluation. The demanding SWAN dataset comprises 32 sequences of point clouds totalling 10,000 frames and containing approximately 0.9 billion points. Sequences 0 to 23 are allocated for training, while sequences 24 to 31 are designated for testing.
Network Configurations: For semantic segmentation on SemanticKITTI, we set the voxel size as 5cm and the maximum number of voxels to 100,000. We use the AdamW optimizer and weight decay as 0.02. The base learning rate is set as 0.004 and the learning rate is scheduled by the Cosine. We adopt rotation, flip, scaling, and transformation to augment training data. On SWAN, we opt not to employ voxelization to reduce point resolution; instead, we directly process the raw point cloud data. We set the maximum number of points to 80,000. We use the AdamW optimizer and set weight decay as 0.04. The base learning rate is set as 0.004 and the learning rate is scheduled by the MultiStepLR. We use the same data augmentation as the ones on SemanticKITTI to preprocess the input data. Our model undergoes training and testing phases with a batch size of 16 and 8 distributed across 4 GPUs.
TABLE V: Semantic segmentation results on the Swan test set.
Methods
mIoU(%)
car
truck
pedestrian
bicycle
motorcycle
@@ Methods
Pointnet++[25] 14.5 31.2 7.3 4.7 9.0 0.0 4.8 0.0 33.9 12.7 59.5 68.4 0.0 13.7 9.0 6.9 15.9 1.0 2.1 11.0 0.0
PointConv[13] 37.3 53.7 20.5 36.5 19.6 5.2 68.7 7.7 61.0 52.5 74.2 77.1 42.3 19.6 38.9 22.1 45.1 26.7 27.3 40.4 6.4
$\psi$-CNN[75] 39.8 48.5 25.2 31.1 22.4 4.2 77.6 7.3 69.0 56.2 73.9 75.6 47.1 23.0 46.1 30.1 57.0 29.1 25.8 38.2 9.4
PolarNet[67] 40.5 78.1 20.3 21.6 4.6 15.3 18.0 8.3 84.0 30.9 91.9 92.7 54.7 33.4 29.5 48.0 60.3 42.7 22.2 42.2 11.8
Cylinder3D[68] 54.9 80.8 30.4 48.7 28.0 6.8 91.7 13.7 85.3 69.0 92.7 92.3 75.2 37.5 72.1 48.8 71.1 46.5 32.5 56.7 18.9
SAT3D[72] 58.2 83.7 45.7 38.7 42.4 11.3 89.5 49.3 85.5 68.2 93.2 92.7 74.9 40.6 77.1 43.3 74.8 42.9 26.7 65.6 17.9

### v2

@@ III-B2 Enhanced Position Encoding
Residual connections are instrumental in training deep neural networks, facilitating gradient flow during backpropagation. Therefore, we combine the Soft Masked Transformer with residual connections to construct a transformer block. As illustrated in Fig. 1 (b), the Soft Masked Transformer Block (SMTB) can be expressed as,
$\vskip-5.69054ptf_{i}$ $={\rm Linear}(f_{in1}),$ (8)
$\mathcal{G}_{i}$ $={\rm SMTransformer}(f_{i}),$
$f_{out1}$ $={\rm Linear}(\mathcal{G}_{i}\oplus f_{i}),$
@@
Figure 2: Comparison of the attention, position encoding in Transformers. (a) The vector attention with position encoding bias in Point Transformer, see Eq.(1). (b) The vector attention with position encoding multiplier in Point Transformer V2, see Eq.(2). (c) The vector attention with soft mask and enhanced position encoding bias in our proposed SMTransformer, see Eq.(3). The different parts are coloured in blue.
@@ III-C Skip Attention-based Up-sampling Block (SAUB)
Conventional upsampling often use simplistic upsampling schemes (e.g., interpolation) that are incapable of effectively recovering the fine-grained geometric details lost during upsampling. To address this problem, we introduce a skip attention-based up-sampling block that combines conventional unpooling with a learnable unit to learn and refine contextual information between features from the encoding and decoding layers across different resolutions, facilitating deep communication between features over various resolution points.
As illustrated in Fig. 1 (c), given the skip feature I $f_{h}\in\mathbb{R}^{{M\times C_{h}}}$ and skip feature II $f_{l}\in\mathbb{R}^{{m\times C_{l}}}$ from the two adjacent encoding layers over the low and high-resolution points, respectively. To build the communication between the different resolution point features, we first balance their feature dimension and point resolution,
$\vskip-5.69054ptf_{mid}={\rm Gridup}\big({\rm Linear}(f_{in2}\oslash f_{l})\big),$ (9)
@@
$\vskip-5.69054pt\mathcal{G}_{i}^{sa}=\sum_{j=1}^{K}\mathcal{A}\big((K_{ij}\ominus Q_{i})\oplus{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\delta(\cdot)}\big)\odot\big(V_{ij}\oplus{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\delta(\cdot)}\big),$ (11)
@@
To compare the differences between SAUB and Conventional upsampling block ( i.e. Transition up[36]), we illustrate their diagrams in Fig. 3. The Skip Attention-Based Upsampling (SAUB) module to facilitate a more intelligent fusion of multi-resolution features. Instead of relying on simple unpooling, SAUB incorporates a learnable attention mechanism that actively learns and refines the contextual information exchanged between encoder and decoder features. This allows it to dynamically adapt to the local point structure, ensuring a detail-preserving and context-aware upsampling process.
Figure 3: Comparison of (a) Transition up and (b) Skip attention-based up-sampling. ‘interpo.’ stands for the interpolation operation. ‘grid up’ stands for the grid-based unpooling. ‘SA’ denotes the skip attention. The different parts are coloured in blue.
Figure 4: (a) Unshared point position encoding: Various transformer blocks (coloured in grey) within the same encoding or decoding layer (coloured in yellow), operating over the same resolution point cloud, require different position information (coloured in blue). (b) Shared point position encoding: Various transformer blocks within the encoding and decoding layers share position information over the same resolution point cloud.
III-D Shared Point Position Encoding
In conventional point transformer networks, several transformer blocks typically utilize different position encoding information within the same encoding or decoding layer across resolution points. We refer to this practice as unshared point position encoding, which would increase the network parameter and training time. Intuitively, points within the same position in the point cloud should have the same position information. Building on this intuition, we propose the shared point position encoding strategy. Under this strategy, transformer blocks within the same encoding or decoding layer share identical position encoding information across resolution points. This approach enhances network robustness and efficiency, particularly in large-scale scene processing. Their comparison is illustrated in Fig. 4.
@@ III-E Network Architecture
We employ a U-Net-like architecture comprising five encoding and decoding layers with skip connections for the semantic segmentation task. The first encoding and decoding layers consist of one MLP and SMTransformer block. The subsequent encoding layers incorporate one Grid Pooling layer [36] followed by several SMTransformer blocks. The number of SMTransformer blocks in the five encoding layers is [1, 2, 2, 6, 2]. The subsequent decoding layers consist of one skip attention-based up-sampling block and one SMTransformer block. We set the feature dimensions as [32, 64, 128, 256, 512] for the 5 encoding and decoding layers. At the network’s end, we append an MLP to predict the final point-wise labels. The network architecture for semantic segmentation is illustrated in Fig. 1 (a).
@@ IV Experimental Results
We evaluate our network on two tasks namely, semantic segmentation of indoor scenes semantic segmentation and outdoor scene semantic segmentation. We also perform detailed ablation studies to demonstrate the effectiveness and robustness of the proposed Soft Mask Transformer Block, Skip Attention-based up-sampling Block and the Shared Point Position Encoding strategy.
@@ IV-A Indoor Semantic Segmentation
Datasets: We evaluate our network on two large-scale indoor scene datasets, namely S3DIS [1] and ScanNetV2 [4]. The S3DIS dataset consists of RGB-D point clouds annotated point-wise with 13 classes. It encompasses 271 rooms from 6 large-scale indoor scenes, totalling 6020 square meters. We utilize 6-dimensional point features, including 3-dimensional normalized colour and 3-dimensional normalized location. For evaluation, we focus more extensively on comparisons using Area 5 as the test set, which is distinct from the other areas and not within the same building.
@@
On ScanNetV2, we set the voxel size as 2cm and the maximum number of voxels to 100,000. We use the Adam optimizer, where the weight decay is set as 0.02. The base learning rate is set as 0.02 and the learning rate is scheduled by the MultiStepLR every 40 epochs. We train and test the model with batch size 24 on 4 GPUs. We adopt random rotation, random scaling, random flip, elastic distortion, chromatic contrast, chromatic translation, chromatic jitter and hue saturation translation to augment training data. The grid size is set as [0.04, 0.12, 0.36, 1.08]cm, and the number of neighbour points is set as 16. During the test, the network uses the test time augmentation, following the Point Transformer V2[36] and Stratified Transformer[16].
TABLE I: Semantic segmentation results on the S3DIS dataset Area-5. We report mean class-wise Intersection over Union (mIoU), mean class-wise accuracy (mAcc), overall accuracy (OA), and parameter (P). w/T denotes the model test with test time augmentation trick. The best result is highlighted in bold. The second best result is underlined
Year Methods w/T mIoU mAcc OA Para.(M)
2024 TMM EPNet[5] $\times$ 72.8 78.3 91.5 14.8
2024 CVPR OneFormer3D[14] $\times$ 72.4
2024 CVPR KPConvX-L[29] $\times$ 72.3 78.4 91.4
2025 AAAI DeepLA-24[41] $\times$ 73.2
2025 IJAEOG U-Next[42] $\times$ 72.7 91.2
2025 IEEE ITJ BeyondPoints[22] $\times$ 73.1 79.1 91.2
2025 AAAI Point Manba[43] $\times$ 79.6
[…]
2021 ICCV Point Transformer[45] $\times$ 70.4 76.5 90.8 7.8
2022 CVPR Stratified Transformer[16] $\times$ 71.4 77.4 90.9 18.8
2022 NIPS Point Transformer V2[36] $\times$ 70.7 76.3 90.9 11.3
2023 NIPS ConDAForMer[6] $\times$ 72.2 78.2 91.2 8.4
2024 CVPR Point Transformer V3[35] $\times$ 72.0 77.8 91.0 46.2
2024 TNNLS Full Point Transformer[8] $\times$ 72.2 78.5 91.5 10.9
SMTransformer(ours) $\times$ 72.6 78.8 91.7 7.8
2021 ICCV Point Transformer[45] $\checkmark$ 70.8 77.0 91.0 7.8
2022 CVPR Stratified Transformer[16] $\checkmark$ 71.9 78.0 91.2 18.8
2022 NIPS Point Transformer V2[36] $\checkmark$ 71.6 77.5 91.0 11.3
2023 NIPS ConDAForMer[6] $\checkmark$ 73.2 78.4 91.2 8.4
2024 CVPR Point Transformer V3[35] $\checkmark$ 73.1 78.3 91.1 46.2
2024 TNNLS Full Point Transformer[8] $\checkmark$ 73.1 78.8 91.7 10.9
SMTransformer(ours) $\checkmark$ 73.6 79.0 91.8 7.8
[…]
$\star$: reproduced the original code by the same workstation.
$\dagger$: denotes the method uses multi-scale semantic information.
All methods do not use multi-dataset joint training
TABLE II: Semantic segmentation results (mIoU) on ScanNetV2 validation set.
Year Methods w/T mIoU(%) Para.(M)
2024 CVPR OneFormer3D[14] $\times$ 76.6
2024 CVPR KPConvX-L[29] $\times$ 75.4
2025 AAAI DeepLA-24[41] $\times$ 74.1
2025 IEEE ITJ BeyondPoints[22] $\times$ 74.6
2025 AAAI Point Manba[43] $\times$ 75.5
[…]
2022 CVPR Stratified Transformer[16] $\times$ 73.0 18.8
2022 NIPS Point Transformer V2[36] $\times$ 74.5 12.8
2023 NIPS ConDAForMer[6] $\times$ 75.2 8.4
2024 CVPR Point Transformer V3[35] $\times$ 76.2 46.2
2024 TNNLS Full Point Transformer[8] $\times$ 74.8 10.9
SMTransformer (ours) $\times$ 75.9 7.8
2022 CVPR Stratified Transformer[16] $\checkmark$ 73.8 18.8
2022 NIPS Point Transformer V2[36] $\checkmark$ 75.0 12.8
2023 NIPS ConDAForMer[6] $\checkmark$ 75.6 8.4
2024 CVPR Point Transformer V3[35] $\checkmark$ 76.6 46.2
2024 TNNLS Full Point Transformer[8] $\checkmark$ 75.6 10.9
SMTransformer (ours) $\checkmark$ 76.2 7.8
$\star$: reproduced the original code by the same workstation.
All methods do not use multi-dataset joint training
TABLE III: Semantic segmentation results on the Swan test set.
@@
motor.
@@
2017 NIPS Pointnet++[26] 14.5 31.2 7.3 4.7 9.0 0.0 4.8 0.0 33.9 12.7 59.5 68.4 0.0 13.7 9.0 6.9 15.9 1.0 2.1 11.0 0.0
2019 CVPR PointConv[34] 37.3 53.7 20.5 36.5 19.6 5.2 68.7 7.7 61.0 52.5 74.2 77.1 42.3 19.6 38.9 22.1 45.1 26.7 27.3 40.4 6.4
2019 CVPR $\psi$-CNN[17] 39.8 48.5 25.2 31.1 22.4 4.2 77.6 7.3 69.0 56.2 73.9 75.6 47.1 23.0 46.1 30.1 57.0 29.1 25.8 38.2 9.4
2021 CVPR PolarNet[44] 40.5 78.1 20.3 21.6 4.6 15.3 18.0 8.3 84.0 30.9 91.9 92.7 54.7 33.4 29.5 48.0 60.3 42.7 22.2 42.2 11.8
2021 CVPR Cylinder3D[46] 54.9 80.8 30.4 48.7 28.0 6.8 91.7 13.7 85.3 69.0 92.7 92.3 75.2 37.5 72.1 48.8 71.1 46.5 32.5 56.7 18.9
2021 ICCV Point Transformer [45] 58.6 84.5 46.2 40.4 42.0 13.3 83.5 50.3 85.2 67.6 93.5 90.3 81.7 51.8 75.1 40.7 72.0 47.0 30.5 61.2 15.3
2022 NIPS Point Transformer V2 [36] 59.7 84.7 47.7 41.2 43.1 14.3 87.5 51.3 86.3 69.2 92.8 91.3 80.2 52.6 76.0 41.5 72.4 48.5 32.2 63.4 17.2
2024 TITS SAT3D[12] 58.2 83.7 45.7 38.7 42.4 11.3 89.5 49.3 85.5 68.2 93.2 92.7 74.9 40.6 77.1 43.3 74.8 42.9 26.7 65.6 17.9
2024 CVPR Point Transformer V3 [35] 61.1 84.7 48.2 45.7 43.3 16.3 88.5 53.2 87.7 69.8 94.8 92.7 81.4 53.3 77.2 45.3 73.8 50.1 33.6 64.1 18.9